RawlsK29:

how do i solve this equation...? -3x = -8x I have absolutely no idea of what to do

RawlsK29:

@Ultrilliam

Ultrilliam:

It cancels itself out, AKA 0

RawlsK29:

Wait... how does it cancel itself out?

Ultrilliam:

You add 8x to both sides, -8x+8x=0 -3x+8x=5x x=0

Ultrilliam:

Vocaloid explain it better pls thx

RawlsK29:

No, i get it. Makes perfect sense. Thank you!(:

Ultrilliam:

Oh, yw! (^_^)